### v1

# Dual Diffusion for Unified Image Generation and Understanding

Abstract
Diffusion models have gained tremendous success in text-to-image generation, yet still lag behind with visual understanding tasks, an area dominated by autoregressive vision-language models. We propose a large-scale and fully end-to-end diffusion model for multi-modal understanding and generation that significantly improves on existing diffusion-based multimodal models, and is the first of its kind to support the full suite of vision-language modeling capabilities. Inspired by the multimodal diffusion transformer (MM-DiT) and recent advances in discrete diffusion language modeling, we leverage a cross-modal maximum likelihood estimation framework that simultaneously trains the conditional likelihoods of both images and text jointly under a single loss function, which is back-propagated through both branches of the diffusion transformer. The resulting model is highly flexible and capable of a wide range of tasks including image generation, captioning, and visual question answering. Our model attained competitive performance compared to recent unified image understanding and generation models, demonstrating the potential of multimodal diffusion modeling as a promising alternative to autoregressive next-token prediction models.
1 Introduction
We are currently in the midst of a multimodal generative modeling revolution. Large scale diffusion models such as Stable Diffusion [19], Dall-E [62], FLUX, and Imagen [64] have become indisputable industry leaders for generating high fidelity images from text descriptions, enabling the accurate modeling and sampling of complex and high dimensional distributions of images given text. Conversely, autoregressive next-token prediction models have achieved groundbreaking performance both in pure text generation and reasoning such as in ChatGPT [1], Gemini [73], and Llama [17] and in visually-grounded text generation with large language models (LLMs), as seen with LLaVA [47] or BLIP-2 [39].
Given these developments, a natural question comes to mind: Can these existing image-to-text (I2T) or text-to-image (T2I) systems be modified to reason with and generate data in the reverse direction? A positive answer would suggest the possibility of producing a fully multimodal model that is able to understand and sample from conditional distributions between modalities in an omni-directional manner. Moreover, unifying these generative frameworks under a single model with shared parameters can confer a multitude of downstream benefits including improved reasoning, simplified implementation, and may be a natural next step towards artificial general intelligence [31, 75].
With autoregressive next-token prediction models, this query has already been answered resoundingly in the affirmative, as evidenced by a multitude of studies [72, 22, 83, 70, 78, 24, 16] demonstrating T2I capabilities of finetuned LLMs. This is in part due to the known next-token generative capability of autoregressive models with visual tokens [82, 74, 40].
On the contrary, with diffusion models there has been surprisingly little evidence of a similar reverse capacity. Until recently, generative diffusion models have struggled with language modeling due to the lack of an empirically performant discrete diffusion process on text tokens, in spite of continued research in this area [4, 15, 42]. At present, multimodal diffusion models either exhibit limited text reasoning capabilities and partial text diffusion [7, 81], which require an autoregressive model such as GPT2 [59] to decode denoised text latents, or emerge as add-ons to pretrained LLMs fine-tuned in conjunction with a diffusion loss [85, 80], and ultimately still rely entirely on next-token prediction for text generation.
We leverage the novel progress in this domain to revisit the above-mentioned question and propose a dual-branch diffusion model based on the multimodal diffusion transformer (MM-DiT) architecture [19], which we modify to output diffusion targets on both modalities of the neural network. We then train our model to perform continuous latent space diffusion on the image branch and discrete masked token diffusion on the text branch. Our novel implementation also allows for controllable infilling in the token space, enabling visual question answering and vision language assistance, which prior diffusion-based models were incapable of. To the best of our knowledge, this is the first end-to-end multimodal diffusion model fully capable of full-featured I2T and T2I generation.
Moreover, we demonstrate the compatibility of our framework with existing diffusion foundation models such as Stable Diffusion 3 (SD3)[19], allowing us to initialize our model with pretrained checkpoints, and reveals remarkably fast adaptation capabilities of the proposed architecture on text generation, producing meaningful text output in under 25B text tokens when initialized with an SD3 checkpoint. Our contributions can be summarized as follows:
  • We introduce a fully end-to-end cross-modal diffusion model that unifies image and text diffusion under a single transformer, which to the best of our knowledge is the first of its kind.
  • We propose a simple, elegant, and easy to implement joint loss function that simultaneously trains the conditional text and image modalities in a unified, end-to-end fashion.
  • We demonstrate performance on an expanded set of multimodal tasks including image generation, visual captioning, and visual question answering using a diffusion-only model, significantly improving on the capabilities and performance of prior multimodal diffusion models.
Modality Task
Image Text Image Image Visual Question
Backbone Backbone Gen Cap. Answering
Versatile Diffusion [81] Diffusion Diff. + AR
Unidiffuser [7] Diffusion Diff. + AR
Show-O [80] Diffusion AR
Transfusion [85] Diffusion AR
Ours Diffusion Diffusion
Table 1: A side-by-side comparison between the backbones and supported features of our work compared to those of existing diffusion-based multimodal methods.
2 Background
[…]
2.1 Continuous Diffusion
Continuous diffusion models operate on continuous vectors by learning to reverse the noise-corruption forward process
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}+\sigma_{t}\boldsymbol{\epsilon},$ (1)
parameterized by time-dependent scalar $\alpha_{t}$ and $\sigma_{t}$, where $\alpha_{t},\sigma_{t}>0$, $\alpha_{t}/\sigma_{t}$ decreases monotonically, and $\boldsymbol{\epsilon}$ is an appropriately selected i.i.d. noise variable. In score-based diffusion models [29, 68], $\alpha_{t},\sigma_{t}$ are determined by a forward stochastic differential equation (SDE) that pushes $\mathbf{x}_{t}$ towards $\mathcal{N}(0,I)$ as tmaps-$t\mapsto\infty$. New samples can be generated by learning the reverse process through estimating the score function [3, 77, 68] $\nabla\log p_{t}(\mathbf{x}_{t})$. Alternatively, from (1), the following ordinary differential equation (ODE) can be derived:
$\dot{\mathbf{x}}_{t}=\mathbf{v}(\mathbf{x}_{t},t),$ (2)
where velocity field $\mathbf{v}(\mathbf{x}_{t},t)=\dot{\alpha_{t}}\mathbf{x}+\dot{\sigma_{t}}\mathbf{\epsilon}$. The ODE in (2) pushes the distribution of $\mathbf{x}_{t}$ from $p_{0}$ to $p_{T}$. To generate new samples, we can use neural networks to approximate the velocity field $\mathbf{v}$ and then integrate ODE (2) backward in time starting from $\mathbf{x}_{T}\sim\mathcal{N}(0,I)$. A common choice of $\alpha_{t},\sigma_{t}$ in flow matching model is $\alpha_{t}=1-t,\sigma=t$ and therefore $\mathbf{v}=\epsilon-\mathbf{x}$, which corresponds to the optimal transport interpolant between two distribution $p_{0}$ and $p_{1}$ [45, 50]. The neural network for regressing the velocity field $\mathbf{v}$ in (2) is trained by optimizing the flow matching loss
$L_{\text{FM}}=\mathbb{E}_{t,q(\mathbf{x}_{t}|\mathbf{x})}||\mathbf{v}_{\theta}(\mathbf{x}_{t},t)-(\epsilon-\mathbf{x})||_{2}^{2}.$ (3)
Recent work such as Stable Diffusion 3 [19] has demonstrated the superiority of flow matching model on text-to-image generation, thus in this work we adopt flow matching objective for modeling the distribution of images.
2.2 Discrete Diffusion
In discrete diffusion, the variate $\mathbf{x}\in\mathcal{X}\times\dots\times\mathcal{X}$ has finite support over the product space of $\mathcal{X}=\{1,\dots,N\}$, where in language models $N$ is the vocabulary size of the token embedding. Generally, there are two ways to approach this modeling task. The first line of works [41, 12, 15, 52, 27] apply a continuous relaxation to the discrete variable and proceed with a continuous reformulation of the framework, allowing the application of the equations in Section 2.1. This greatly simplifies the diffusion modeling itself, but introduces a significant source of error in the mapping between discrete and relaxed continuous states. Conversely, the diffusion process is extended to the discrete token space [4, 54, 51, 65], which removes the need for the aforementioned mapping via a specialized discrete diffusion formulation. In our work, we will focus on this latter perspective given its empirical potential as validated in recent work [51, 65, 23, 66]. Leveraging continuous-time Markov chain (CTMC) theory, the marginal distributions $p_{t}$ can be described by a family of linear ordinary differential equations
$\frac{dp_{t}}{dt}=Q_{t}p_{t},$ (4)
where $p_{0}\approx p_{\text{data}}$ and $p_{1}=p_{\text{stationary}}$, and $Q_{t}$ is a time-dependent sequence of transition matrices that provides a mapping between the two distributions. We consider the case of absorbing state (i.e., masked) diffusion that are shown to work well on text modeling [51, 65, 66]. This formulation induces the posterior ($0<s<t$)
$q(\mathbf{x}_{s}|\mathbf{x}_{t},\mathbf{x})=\begin{cases}\text{Cat}(\mathbf{x}_{s}|\mathbf{x}_{t})&\mathbf{x}_{t}\neq\mathbf{m}\\ \text{Cat}(\mathbf{x}_{s}|\frac{(1-\alpha_{s})\mathbf{m}+(\alpha_{s}-\alpha_{t})\mathbf{x}}{1-\alpha_{t}})&\text{o.w.}\\ \end{cases}$ (5)
where clean data $\mathbf{x}$ is a discrete variable (one-hot vector) with $N$ categories, with the marginal
$q(\mathbf{x}_{t}|\mathbf{x})=\text{Cat}[\mathbf{x}_{t}|\alpha_{t}\mathbf{x}+(1-\alpha_{t})\mathbf{m}],$ (6)
where $\text{Cat}(\cdot|\boldsymbol{\pi})$ denotes the categorical distribution over different classes with probability $\boldsymbol{\pi}$, and $\mathbf{m}$ denotes the mask absorbing state.
To reverse this process, one may either model the density ratio $s_{\theta}(\mathbf{x})_{\mathbf{y}}\approx\frac{p_{t}(\mathbf{x})}{p_{t}(\mathbf{y})}$ given two sequences $\mathbf{x},\mathbf{y}\in\mathcal{X}\times\dots\times\mathcal{X}$ as in Uniform Score-Entropy Discrete Diffusion (SEDD) [51], or the denoised variate $\mathbf{x}_{\theta}(\mathbf{x}_{t},\alpha_{t})\approx\mathbf{x}$ directly as in masked diffusion model [65, 66]. In the former, the modeled density ratios induce a specialized reverse transition matrix $\bar{Q}_{t}$ that can be leveraged in Eq. 4. In the latter, $\mathbf{x}_{\theta}$ can be directly substituted for $\mathbf{x}$ in Eq. 5. In this work, we follow Sahoo et al. [65] that enforces zero-probability on the mask state $\mathbf{m}$ and keeps all un-masked state unchanged during reverse sampling. This induce a simplified (negative) variational lower bound under the continuous time limit
[…]
In practice, we can use Monte-Carlo sampling to approximate and evaluate the loss function in (7). Following Sahoo et al. [65], we use a log-linear schedule: $\alpha_{t}=1-t$.
Figure 1: Our proposed model, the Dual Diffusion Transformer (D-DiT) that simultaneously models image and text distributions via a joint denoising diffusion training loss. a) An overview of the model architecture. The gray blocks (T5 encoder, image autoencoder) are kept fixed throughout training and inference. b) During training for (image-conditioned) text denoising, the text input is randomly masked while the image is noise-free. c) During training for text-conditioned image denoising, the image is randomly noised while the text is noise-free.
3 Method
We propose an end-to-end multi-modal diffusion model named Dual Diffusion Transformer (D-DiT) with a unified backbone that jointly models image and text distribution. More specifically, given image $\mathbf{x}^{(\text{img})}$ and text $\mathbf{x}^{(\text{txt})}$, we are interested at modeling the conditional distribution $p(\mathbf{x}^{(\text{img})}|\mathbf{x}^{(\text{txt})})$ and $p(\mathbf{x}^{(\text{txt})}|\mathbf{x}^{(\text{img})})$. The former is usually referred to as text-to-image generation and the latter is closely related to various image understanding tasks such as captioning and visual question answering.
3.1 Architecture
Inspired by the MM-DiT in SD3 [19], our proposed D-DiT is a Transformer-based model comprising two branches - one for processing image tokens and another for processing text tokens. The image and text tokens attend with each other in every attention layer. In D-DiT, the output of the image branch is the prediction of velocity defined in (2) with text conditioning, while the output for the text branch is the $\mathbf{x}^{(\text{txt})}$ prediction with image conditioning. The scalar timestep embedding modulates every layer’s feature map via AdaLN (adaptive layernorm) [56]. We only input the timestep $t$ to the model during image generation, as $\mathbf{x}^{(\text{txt})}_{t}$ has already implicitly carried the information of signal-to-noise ratio (the number of mask tokens in the sequence). In addition, we add a text encoder with bi-directional attention on top of the text branch of the diffusion model. Despite an asymmetric design of the model is not strictly required, having a text encoder on top of a DiT model allows us to easily adapt many existing text-to-to-image model such as SD3 and FLUX as pretrained backone for our D-DiT model. Note that the text encoder should not have causal mask as this will violate the masked diffusion process.
To reduce the computational cost associated with high-resolution images, we follow prior works on latent-space (image) diffusion [63] which compresses images from raw pixel space to a spatially compressed latent space obtained from a variational-autoencoder (VAE) trained with discriminator loss [18] and KL-divergence regularization [34].
3.2 Training
We propose a joint training objective for image-text joint modeling, which is essentially a joint denoising target that combines continuous and discrete diffusion. Formally, we use flow matching introduced in Section 2.1 to learn the conditional distribution of images and masked diffusion introduced in Section 2.2 to learn the conditional distribution of texts. During training, corrupted samples $\mathbf{x}_{t^{(\text{img})}},\mathbf{x}_{t^{(\text{txt})}}$  [^1] For better readability superscript $(\text{txt}),(\text{img})$ are omitted. are sampled from the corresponding forward corruption processes $q(\mathbf{x}_{t}|\mathbf{x})$ defined in (1) and (6) respectively and then we can calculate the diffusion loss for each modality
$L_{\text{image}}$ (8)
$=\mathbb{E}_{t,q^{(\text{img})}}\left|\left|\mathbf{v}_{\theta}\left(\mathbf{x}^{(\text{img})}_{t},t,\mathbf{x}^{(\text{txt})}\right)-(\boldsymbol{\epsilon}-\mathbf{x}^{(\text{img})})\right|\right|_{2}^{2},$
$L_{\text{text}}$
$=\mathbb{E}_{q^{(\text{txt})}}\left[-\frac{1}{K}\sum\nolimits_{i=1}^{K}\log[\mathbf{x}_{\theta}(\mathbf{x}_{t_{i}}^{(\text{txt})},\mathbf{x}^{(\text{img})})\cdot\mathbf{x}]/t_{i}\right],$
where we use antithetic sampling for text diffusion timesteps $t_{i}$ by discretizing $(\delta,1]$ into $K$ points uniformly with $\delta$ being a small number to avoid numerical instability, for image diffusion we sample $t$ from the log-normal distribution. We do not corrupt the conditioning samples during training and as such the image diffusion timestep is always set to zero when predicting text distribution.
To sum up, the overall dual modality training loss is a simple weighted combination of the above single modality diffusion loss:
$L_{\text{dual}}=L_{\text{image}}+\lambda_{\text{text}}L_{\text{text}},$ (9)
with $\lambda_{\text{text}}$ being a hyperparameter.
3.3 Inference
We introduce three types of sampling-based inference which can be used for different vision-language tasks, which we detail below.
Text-to-image Generation
To perform text-guided image generation, i.e. $\mathbf{x}\sim p(\mathbf{x}^{(\text{img})}|\mathbf{x}^{(\text{txt})})$, we use the commonly adopted classifier-free guidance (CFG) technique [28] to sample from the conditional distribution $p(\mathbf{x}^{(\text{img})}_{t}|\mathbf{x}^{(\text{txt})})$, which amounts to a re-weighting of the velocity prediction
$\tilde{v}_{t}=s\mathbf{v}_{\theta}\left(\mathbf{x}^{(\text{img})}_{t},t,\mathbf{x}^{(\text{txt})}\right)+(1-s)\mathbf{v}_{\theta}\left(\mathbf{x}^{(\text{img})}_{t},t,\emptyset\right),$ (10)
where $s$ is a hyperparameter that controls the scale of guidance and $\emptyset$ is a suitable null embedding (e.g. the embedding of an empty text) .
Image-to-text Generation
To sample images from the conditional distribution, we can use ancestral sampling to draw from the posterior distribution $q(\mathbf{x}_{s}|\mathbf{x}_{t},\mathbf{x})$ in (5) by plugging in prediction $\mathbf{x}\approx\mathbf{x}_{\theta}(\mathbf{x}_{t}^{(\text{txt})},\mathbf{x}^{(\text{img})};t=0)$.
Image-to-text In-filling
In certain tasks, both text conditioning information and image conditioning information are available, such as in a visual question answering task where an image and an associated question are provided. For such cases, we would like to sample $\mathbf{x}\sim p(\mathbf{x}^{(\text{answer})}|\mathbf{x}^{(\text{img})},\mathbf{x}^{(\text{question})})$.
To perform this task, we initialize the diffusion prior of the question with masked tokens and leverage the robust text in-filling capabilities of the text diffusion model to complete the sequence by sampling from the conditional distribution. The text question tokens are kept fixed throughout sampling (Figure 2).
Figure 2: Text masking during both training and sampling under the image captioning (above) and visual question answering (below) tasks with our proposed model.
4 Experiments
Figure 3: Text-to-image samples generated from the model. We draw images from the reverse diffusion process via the Euler solver with $T=28$ diffusion steps.
4.1 Experimental Setup
Implementation details
We implement our proposed framework based on the open-sourced SD3-medium model [19] [^2] https://huggingface.co/stabilityai/stable-diffusion-3-medium. We initialize the model weight of the DiT from the pretrained checkpoint and add a linear head on top of the text branch to do text denoising. Following SD3, we adopt the existing T5 encoder/tokenizer [61], and SD3’s image VAE, whose weights remain unchanged throughout all the experiments (except for the mask token embedding in T5). We remove the CLIP text encoders in the SD3 model due to its causal attention mask and for a simplified model structure. We use the special token <extra_id0> in T5’s vocabulary to represent the mask token in masked diffusion, as this token is used to mark the masked token in the mask pretraining process of original T5 model. In this way, we find the model can generate text reasonably well even without updating the weight of this token embedding. To further reduce the domain gap, we unfreeze the token embedding of <extra_id0> during the second stage of the training.
Different from multi-modal models that are built upon language models, our model has never been trained on text-only generation. In the preliminary experiments, we found that adding text-only target (i.e. un-conditional text generation) to the model does not influence its captioning performance significantly. An interesting future direction can be extending the proposed framework to model the marginal distribution of each modality.
Datasets
We train the model with three stages on publicly available datasets. The total image-text pairs we have used is roughly 40M. We list the detail of the dataset and training setup for each stage below, where all the training stages use the joint diffusion loss defined in (8).
  1. 1.
    Dual diffusion pretraining. The original SD3 model was only trained on ambient image-text pairs, and not solely on text data itself. To adapt D-DiT to text generation tasks, we train it on the joint diffusion loss for 60K iterations with a batch size of 512. The maximum text token length is truncated to 64 and we use image resolution of 256. The dataset used in this stage is re-captioned Datacomp-1b [21, 43] (the model has only seen around 30M images in this stage, which is less than $3\%$ of the total images in the dataset).
  2. 2.
    Continued pretraining on higher quality data. We then unfreeze the masked token embedding in T5 and train the model for 200k iterations on a image understanding dataset with rich textual description, which consists of the pretraining dataset from ShareGPT4V[11] (1.3M images) and OpenImages (1.9M subset with object detection annotations) [35] re-captioned by ShareCaptioner [^3] https://huggingface.co/Lin-Chen/ShareCaptioner. The text token length is set to 256 and image resolution is 256, with batch size of 512. Finetuning the mask token embedding reduces the domain gap as T5 encoder has not seen sequence filled with high percentage of mask token during its pretraining.
    However, as updating the mask token embedding requires backpropagating through the heavy T5 encoder, we freeze the mask embedding after this round of training on the image understanding dataset. We observe that the $\ell 2$ difference between the mask token embedding from different training iterations does not change much after 100k iterations.
    Moving forward, we conduct an optional high resolution model finetuning on the aforementioned image understanding dataset together with a newly added image generation dataset. Concretely, we collect a higher quality dataset with 10M images (9M re-captioned LAION-1024 and 1M midjourney images [^4] https://huggingface.co/datasets/CaptionEmporium/midjourney-niji-1m-llavanext). In this training stage, the image diffusion loss is calculated on the high quality image dataset whereas the text diffusion loss is calculated on the understanding dataset. We finetune the model for 80k iterations, with image resolution 512, text token length 256, and a batch size of 768. Only our 512$\times$512 model variant requires this training stage.
  3. 3.
    Visual instruction tuning. Finally, we finetune our model on an medley of instruction-tuning datasets to promote joint text-image conditioned text generation. We combine the LLaVA-Pretrain558K and LLaVA-v1.5-mix-665K visual instruction tuning datasets with the training splits for TextVQA and VizWiz and train for 25k iterations. Following the convention in LLaVA-1.5, the model is trained to distinguish between long-form answers and short answers, multiple choice answers, or captions via task-specific instruction prompts that come after the question such as ”Answer the question using a single word or phrase,” ”Answer with the option’s letter from the given choices directly,” or ”Describe the image concisely.”
4.2 Multi-modal Understanding
Existing multi-modal diffusion models such as UniDiffuser [7] and Versatile Diffusion [81] performed text diffusion in a CLIP latent space, which hampered their ability to perform text completion, a necessary feature for general question answering and conversation-based tasks. This is no longer a limitation with our proposed D-DiT due to its discrete masked diffusion branch, allowing us to leave question tokens unmasked throughout sampling. We are thus able to evaluate our fine-tuned model on a full suite of image-to-text generation tasks, including image captioning and visual question answering benchmarks, as well as long-form visual assistance responses.
We first evaluate the visual understanding capabilities of D-DiT via the academic question answering benchmarks VQAv2 [26], VizWiz [8], OKVQA [53], GQA [30], POPE [44], as well as MME [20]. Due to the short-form nature of the questions, we perform sampling with 16 diffusion steps, and compare against a selection of multi-modal models, including I2T only and I2T + T2I models. Our results are summarized in Table 3. We note that our D-DiT as the only diffusion-only multi-modal model capable of visual question answering tasks, already boosts performance that is competitive with recent I2T + T2I models. Our model at 512 resolution outperforms Show-O on MME, GQA, and POPE, approaching performances of auto-regressive VLMs such as QWEN-VL and BLIP-2.
Next, we provide qualitative examples of the D-DiT, providing images and gauging the model’s visual language assistance capabilities via image-related queries. Given the longer format of the responses, we sample D-DiT responses with 256 diffusion steps. Our model provides answers to human queries in a manner that suggests a fine-grained multi-modal understanding of the image and text conditioning (Figure 4).
4.3 Text-to-image Generation
Besides the image-conditioned text generation, we also test model’s text-to-image generation capability. Following previous works, we evaluate our 512$\times$512 model after the second training stage on the GenEval benchmark, which measures model’s prompt following capability [25]. We follow the default setting in the open-sourced SD3 checkpoint where we use a Euler solver with 28 sampling steps and a CFG scale of $7.0$. We observe that the joint diffusion training does not cause catastrophic forgetting on the model, the fine-tuned D-DiT preserves the performance of original SD3 model and slightly improves on some of metrics such as colors after joint training. Qualitative evaluation samples are shown in Figure 3, where we observe the ability to generate highly aesthetic images is preserved.
Model params (B) Overall Objects Counting Colors Position Color attribution
Single Two
PixArt-$\alpha$ [10] 0.6 0.48 0.98 0.50 0.44 0.80 0.08 0.07
SD V2.1 0.9 0.50 0.98 0.51 0.44 0.85 0.07 0.17
DALL-E 2 [62] 6.5 0.52 0.94 0.66 0.49 0.77 0.10 0.19
SDXL [57] 0.9 0.55 0.98 0.74 0.39 0.85 0.15 0.23
DALL-E 3 - 0.67 0.96 0.87 0.47 0.83 0.43 0.45
CoDI [71] - 0.31 0.89 0.16 0.16 0.65 0.02 0.01
LWM [48] 7 0.47 0.93 0.41 0.46 0.79 0.09 0.15
SEED-X [24] 17 0.49 0.97 0.58 0.26 0.80 0.19 0.14
Chameleon [72] 7 0.39 - - - - - -
Show-O [80] 1.3 0.68 0.98 0.80 0.66 0.84 0.31 0.50
Transfusion [85] 8 0.67 - - - - - -
SD3 [19] 2 0.62 0.98 0.74 0.63 0.67 0.34 0.36
D-DiT (ours) 2 0.65 0.97 0.80 0.54 0.76 0.32 0.50
Table 2: Evaluation of text-to-image generation performance on Geneval [25]. params denote the number of trainable parameters.
Figure 4: Multi-modal dialogue examples generated from our model. To our knowledge, D-DiT is the first diffusion-based multimodal model capable of instruction-based vision and language conversation.
Model Params Text Image MS-COCO VQAv2 VizWiz OKVQA MME GQA POPE
# trainable Backbone Backbone CIDEr $\uparrow$ Acc. $\uparrow$ Acc. $\uparrow$ Acc. $\uparrow$ Acc. $\uparrow$ Acc. $\uparrow$ Acc. $\uparrow$
LLaVA-1.5 [46] 13B AR - - 81.8 57.5 - 1500.1 64.7 86.4
BLIP-2 [39] 13B AR - - 65.0 19.6 - 1293.8 41.0 85.5
IDEFICS [36] 9B AR - - 50.9 - - - - -
QWEN-VL [6] 7B AR - - 78.2 38.9 - 1487.5 57.5 -
OpenFlamingo [5] 9B AR - 65.5 43.5 - - - - -
Flamingo [2] 9B AR - 79.4 51.8 28.8 44.7 - - -
CM3Leon [83] 7B AR AR 61.6 47.6 37.6 23.8 - - -
Chameleon [72] 7B AR AR 18.0 - - - - - -
LWM [49] 7B AR AR - 55.8 11.6 - - 44.8 75.2
Show-O (256$\times$256) [80] 1.3B AR Diffusion - 64.7 - - 1014.9 54.2 76.2
Show-O (512$\times$512) [80] 1.3B AR Diffusion - 69.4 - - 1097.2 58.0 80.0
Transfusion [85] 7B AR Diffusion 29.0 - - - - - -
D-DiT (Ours, 256$\times$256) 2B Diffusion Diffusion - 59.5 19.4 28.5 897.5 55.1 79.2
D-DiT (Ours, 512$\times$512) 2B Diffusion Diffusion 56.2 60.1 29.9 25.3 1124.7 59.2 84.0
Table 3: Comparison of our D-DiT against related work on visual question answering benchmarks. VLMs that focus on text-generation remain superior to unified understanding and generation models, however our models compare favorably with the latter category.
4.4 Ablation Studies
As text-to-image diffusion models are trained on a large number of text-image pairs, one may raise the question, whether the representation learned throughout this process can be transferred to multi-modal understanding tasks? To answer this question, we perform an ablation study on the internal representation of a text-to-image diffusion model. We adapt several models into an image captioning model, including SD3, CLIP ViT L/14, and our D-DiT model. Among different internal layers in SD3, we find the feature from the 18th layer tend to perform the best in our preliminary experiment so it is used as the output feature. We add an GPT2 text decoder to the features extracted from SD3 and CLIP, and directly use D-DiT’s text output as results. We train all models with a mixture of recaptioned Datacomp, recaptioned OpenImages and captioning data from ShareGPT4V [11]. Concretely, we evaluate the quality of the captions generated from the models by asking GPT4 to do visual question answering according to the generated captions. The accuracy is listed in Table 4.
Similar to the trend observed in [75], directly using diffusion features as the prefix of a language decoder yield worse performance compared to language-supervised vision models like CLIP ViT [60]. Unfreezing the parameters of the diffusion backbone slightly improve the performance, but it still cannot match the CLIP encoder. This suggests that the representation from image diffusion models is not directly transferrable to the text embedding space where the decoder-only language model operates on. Instead of leveraging a separate language decoder, we use the text branch in the MM-DiT architecture to directly model the conditional text distribution, which notably boost the performance. This uncovers an intriguing property of MM-DiT model or potentially other bi-directional Transformers, that these models are good representation learners for estimating the likelihood of multi-modality data distribution.
We also conduct an ablation study with respect to the number of text diffusion sampling steps and study its influence on VQA accuracy with VQAv2 and captioning quality on COCO dataset. For VQAV2, which involves short text answers, good accuracy can be achieved with relatively few sampling steps. For the captioning task on MS-COCO, performance improves as the number of sampling steps increases, mirroring the trend observed by Sahoo et al. [65], where additional sampling steps lead to reduced perplexity.
Vision Encoder Language Decoder VQAv2 (val)
0-shot 32-shot
SD3 feature (frozen) GPT 2 42.3 46.9
SD3 feature (trainable) GPT 2 45.1 50.2
CLIP ViT L/14 (frozen) GPT 2 50.6 54.8
UniDiffuser [7] GPT 2* 46.7 49.4
D-DiT (ours) - 55.0 60.3
Table 4: Comparison between different vision encoder and proposed model. *The GPT 2 decoder of UniDiffuser is finetuned on text reconstruction and kept frozen afterwards.
Task $T=4$ $8$ $16$ $32$ $64$ $128$
VQAV2 (acc.) 58.8 58.0 59.3 60.5 60.0 59.6
MS-COCO (CIDEr) 20.2 35.3 46.5 51.3 56.2 54.5
Table 5: An ablation study on the effect of sampling steps $T$ on discrete text diffusion performance in terms of COCO Captioning CIDEr score and VQAV2 subset’s question-answering accuracy.
5 Related Works
5.1 Diffusion Models
Diffusion models [67, 29, 68] generate data by gradually converting noise into signal via a reverse diffusion process. They are the de facto standard for image generation [14, 32] and likelihood modeling [33, 55, 69, 38]. Conditional diffusion models [28] have also been shown to be powerful interfaces bridging text and images, particularly for their ability to generate highly realistic and aesthetic images from textual descriptions [63, 64, 57, 22, 10, 62]. Their exceptional performance in the image domain has also inspired numerous extensions to the language generation [41, 12, 15, 52, 27, 51, 65, 23, 66], and is an attractive alternative as its sampling is not constrained by a specified token generation order and the attention mechanism does not need to be uni-directional.
5.2 Vision Language Models
The success of large language models (LLMs) [9, 76] and vision-language pretraining [60] has given rise to a series of multi-modal language models. The visual signal is projected to the text embedding space via vision encoders supervised by text labels [60, 84] and then connected to a pretrained language models through further instruction tuning [47, 86, 13, 2, 75]. While these models have shown promising capabilities in image understanding and few-shot generalization, their predictive targets are inherently language-centric, limiting their ability to model the image distributions directly.
5.3 Multimodal Text and Image Generative Models
Rather than simply connecting visual encoders to language models, recently there has been an active line of inquiry focused on exploring a unified generative model for joint vision and language generation. Inspired by autoregressive language models, many of the unified multi-modal generative models extend the next-token prediction to both image and text tokens [70, 78, 24, 16, 72]. More recently, Transfusion [85] and Show-O [80] demonstrate that bi-directional image diffusion can be integrated with autoregressive text prediction in the same framework. On the other hand, Versatile Diffusion [81] and Uni-diffuser [7] explore applying a continuous diffusion process to text and image modalities, where text generation is broken into two stages - first, continuous diffusion is used to generate latent embeddings which are then decoded into text by another LLM (e.g. GPT2 [58]). While these works hint at the potential of diffusion models as efficient multi-modal models, their text generation capability is restricted to simple tasks like generating short captions from images.
Conclusion and Discussion
In this work, we introduced an end-to-end multi-modal diffusion model that bridges the gap between text and image diffusion by enabling both text-to-image (T2I) and image-to-text (I2T) tasks through a unified diffusion model. We demonstrated that a bi-directional transformer trained with a joint diffusion target is an effective multi-modal learner capable of competing with the autoregressive models that have long dominated the field. Additionally, the bi-directional attention mechanism is equivariant to the order of input tokens, enabling the prediction of conditional distributions without requiring a specific arrangement of different modalities or special handling of the attention mask.
Limitations
While discrete diffusion offers the advantage of being agnostic to sequential order and is compatible with bi-directional Transformers, its current implementation requires the sequence length to be preset before sampling. A promising future direction would be to extend the sampling scheme to allow for more flexibility, enabling dynamic sequence lengths during the sampling process. In addition, while we show that our proposed dual diffusion model can perform instruction tuning, its instruction-following capabilities still marginally lag behind those of state-of-the-art autoregressive models.
Supplementary Material
6 Training Details
Hyperparam. Dual pretrain Continued pretrain Instruct. tuning
Mask emb. High res.
Gradient steps 60k 200k 80k 25k
Batch size 512 512 768 512
LR 5e-5 3e-5 3e-5 3e-5
Scheduler Constant LR with warmup
Warmup iters 5000 1000 1000 1000
Weight decay 1e-2
Text loss weight 0.2 0.3
Table 6: Training hyperparameters for D-DiT. Text loss weight denotes the $\lambda$ in (8).
We provide the detailed hyperparameter setting for different training stages in the Table 6. During all the training stages, we use AdamW optimizer with default hyperparameters ($\beta_{1}=0.9,\beta_{2}=0.999$). Mixed precision training (bf16) and fully-sharded data parallel (with gradient and optimizer state sharded) are used for model training.
7 Further Results
Model Backbone Params. (B) FID $\downarrow$
SD-XL [57] Diff. 0.9 9.55
PixArt-$\alpha$ [10] Diff. 0.6 6.14
Playground v2.5 Diff. - 4.48
Show-O [80] Discrete Diff. 1.3 15.18
LWM [48] AR 7 17.77
VILA-U [79] AR 7 7.69
SD3 [19] Diff. 2 16.45
D-DiT Diff. 2 15.16
Table 7: Comparison with other models on MJHQ-30K evaluation benchmark at $512\times 512$ resolution.
We evaluate the aesthetic quality of generated images from our proposed D-DiT against those of the original SD3 model and a selection of existing text-to-image (T2I) and multi-modal works. We measure Frechet Inception Distance (FID) with respect to a collection highly aesthetic generated images, known as the MJHQ-30K benchmark proposed by [37]. As shown in Table 7, we observe an improvement in FID after joint diffusion training, and favorable comparison against multi-modal models of similar size.
We provide an illustrative example of masked diffusion in Figure 5 for the visual question answering task, where the token generation process is visualized over diffusion time. Over the course of sampling, the answer tokens are gradually denoised from the masked state via absorbing state reverse diffusion. The question tokens are always left unmasked throughout the entire process.
We also include a qualitative comparison in captioning performance compared to UniDiffuser [7], another diffusion-based multi-modal model, in Figure 6, where we demonstrate an improvement in the ability to capture fine-grained details of the image in a longer caption format. Finally, we provide further uncurated text-to-image (T2I) generation results in Figures 7, 8, 9, and 10. Overall, these results further demonstrate the multi-faceted performance of our proposed dual-branch diffusion-based multi-modal model.
Figure 5: Illustrative example of visual question answering with mask diffusion.
Figure 6: Comparison of captions generated by D-DiT and UniDiffuser[7]. The prompt to D-DiT is ”Provide a brief description of the given image.”
Figure 7: Additional text-to-image samples generated from the model.
Figure 8: Additional text-to-image samples generated from the model.
Figure 9: Additional text-to-image samples generated from the model.
Figure 10: Additional text-to-image samples generated from the model.